Susan drives 240 miles in four and a half hours. What is her average rate during the drive? Round your answer to the nearest tenth.
Answer by ewatrrr(24785)


Hi,
240 miles in four and a half hours
d= r*t Or r = d/t
average rate
240mi/4.5hr = 53.3 mi/hr (to the nearest tenth)